OpenStudy (anonymous):

Joe bought a new fish aquarium. It holds 10 gallons of water and Joe wants to fill it three quarters of the way up. Unfortunately, Joe does not have a bucket to carry water to the aquarium. In fact the only thing he has is an eight ounce glass. He can fill the glass and dump it into the aquarium in 45 seconds. Joe starts filling the aquarium at noon, if he can keep up this constant rate at what time will he finish?

OpenStudy (johnweldon1993):

Oh geez Joe has some work to do.. First off....How many ounces are in 1 gallon?

OpenStudy (cwrw238):

lol - a lot of work I guess you are talking about an American gallon. I think thats different from a British gallon. right?

OpenStudy (anonymous):

@johnweldon1993 1 gallon = 128 ounces @cwrw238 yes American gallon

OpenStudy (johnweldon1993):

Right...so 128 ounces in 1 gallon...so in 10 gallons there will be 128 times 10 right? So 1280 ounces in the entire fishtank Now luckily...he doesnt need to do it ounce for ounce....he has an 8 ounce cup... So how many 8 ounce cups are in that 1280 ounces? \[\large \frac{1280}{8}= ?\]

OpenStudy (anonymous):

@johnweldon1993 160 eight-ounce cups in 1280 ounces

OpenStudy (johnweldon1993):

Right... So what does that mean? That means..in order to fill up the entire aquarium...he needs to make 160 trips....of 8 ounces of water to fill up that tank (would have me investing in a bucket)

OpenStudy (johnweldon1993):

So now...if he makes 160 trips....and each trip takes 45 seconds....how many seconds is that? \[\large 160 \times 45\]

OpenStudy (anonymous):

7200 seconds @johnweldon1993

OpenStudy (johnweldon1993):

Right...which is how many minutes? 60 seconds in 1 minute so \[\large 7200 / 60 =?\]

OpenStudy (anonymous):

@johnweldon1993 120

OpenStudy (johnweldon1993):

Right... so...if Joe started at Noon.... And it takes him 120 minutes to do this whole ordeal....what time does he end? 12:00 + 120 minutes... Just so happens....120 minutes / 60 minutes in 1 hour = 2 hours... So if it takes him 2 hours to complete this..the time will be ...?

OpenStudy (anonymous):

2:00 @johnweldon1993

OpenStudy (johnweldon1993):

There you go man!

OpenStudy (anonymous):

Thanks so much! You made it so simple! @johnweldon1993

OpenStudy (johnweldon1993):

No problem!